OpenStudy (rianiscool):

need help asap will give a medal Which is an example of an algebraic expression? A. 4 + c > 1 B. b > –2 C. 12 = a + 6 D. m – 3

OpenStudy (igreen):

Expressions have no equal signs or inequality signs..so what's the answer? @rianiscool

OpenStudy (igreen):

Inequality Signs Are: \(<, >, \le, \ge\)

OpenStudy (rianiscool):

?

OpenStudy (igreen):

Algebraic expressions don't have equal signs(\(=\)) or inequality signs(\(<, >, \le, \ge\)).

OpenStudy (igreen):

So what's the algebraic expression here?

OpenStudy (perl):

D

OpenStudy (rianiscool):

that is my ?

OpenStudy (igreen):

@perl Don't give out answers.

OpenStudy (perl):

the first three are inequalities

OpenStudy (perl):

oh sorry

OpenStudy (igreen):

The first three aren't inequalities, either.

OpenStudy (perl):

woops, first 2 i meant

OpenStudy (perl):

first 2 inequalities, third is an equation. fourth is a polynomial (algebraic expression

OpenStudy (rianiscool):

so it would be b

OpenStudy (igreen):

Yeah. The first two have the greater than sign, which is an inequality. The third one has an equal sign. The fourth one is an expression.

OpenStudy (yousir):

y did u block me

OpenStudy (igreen):

It's not going to be B..look over what we said. @ria

OpenStudy (igreen):

@rianiscool

OpenStudy (igreen):

Who are you talking to? @YouSir

OpenStudy (rianiscool):

ok

OpenStudy (igreen):

Which is an example of an ALGEBRAIC EXPRESSION?

OpenStudy (yousir):

pearl

OpenStudy (igreen):

Oh.

OpenStudy (rianiscool):

i am bad at math

OpenStudy (rianiscool):

@iGreen

OpenStudy (igreen):

We're looking for an ALGEBRAIC EXPRESSION. The first two options have the greater than sign, which is an INEQUALITY. The third option has an equal sign, which is an EQUATION. The fourth option is an EXPRESSION. @rianiscool So what's the answer?

OpenStudy (rianiscool):

D

OpenStudy (igreen):

Yep.

OpenStudy (rianiscool):

ok thank you @igeen but who gets the medal

OpenStudy (rianiscool):

@iGreen

OpenStudy (igreen):

Whoever you think deserves it the most.

OpenStudy (rianiscool):

]\ ok thank you